OpenStudy (anonymous):

f-32/9=c/5 solve for f

OpenStudy (anonymous):

????????????

OpenStudy (texaschic101):

f - 32/9 = c/5 -- multiply by the common denominator, which is 45 45f - 160 = 9c -- add 160 to both sides 45f = 9c + 160 -- divide both sides by 45 f = (9c + 160) / 45 OR f = 1/5c + 32/9